OpenStudy (luigi0210):

Factor:

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (luigi0210):

\(\LARGE \color{green}{ x^2+4x+5} \)

OpenStudy (alexandervonhumboldt2):

well usec ompleting the square method

OpenStudy (alexandervonhumboldt2):

x^2 + 2*2x + 4 +5 -4 (x+2)^2 -1

OpenStudy (alexandervonhumboldt2):

1^2=1 difference of squares: (x+2-1)*(x+2+1)

OpenStudy (alexandervonhumboldt2):

(x+2-1)*(x+2+1)=??? simp,e math

OpenStudy (luigi0210):

Just out of curiosity, did you get \(\Large 2 \pm 2i \)?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

oh wait

OpenStudy (alexandervonhumboldt2):

you mean as a root ?

OpenStudy (luigi0210):

Whoops, *1i

OpenStudy (alexandervonhumboldt2):

wait the factored form will be with no i

OpenStudy (luigi0210):

It would if you checked it right :P

OpenStudy (luigi0210):

\(\LARGE \frac{-4 \pm \sqrt{4^2-20 }}{2} \) ^

OpenStudy (alexandervonhumboldt2):

woops i made amistake somewhere let me check

OpenStudy (luigi0210):

Haha, it's alright, just wanted to see if you would make the same mistake as others ;P Thanks for your time Alexander :3

OpenStudy (alexandervonhumboldt2):

x^2+4x+5 x^2+2*2x+4+5-4 (x+2)^2+1 yeah i mistaked in sing

OpenStudy (alexandervonhumboldt2):

so yeah complelting the square methos is not suitable here you need to use quadratic formula

OpenStudy (alexandervonhumboldt2):

use quadratic formula to find root well yo u already found them